### v1

# Tangling-Untangling Cycle for Efficient Learning

Abstract
The conventional wisdom of manifold learning is based on nonlinear dimensionality reduction techniques such as IsoMAP and locally linear embedding (LLE). We challenge this paradigm by exploiting the blessing of dimensionality. Our intuition is simple: it is easier to untangle a low-dimensional manifold in a higher-dimensional space due to its vastness, as guaranteed by Whitney embedding theorem. A new insight brought by this work is to introduce class labels as the context variables in the lifted higher-dimensional space (so supervised learning becomes unsupervised learning). We rigorously show that manifold untangling leads to linearly separable classifiers in the lifted space. To correct the inevitable overfitting, we consider the dual process of manifold untangling - tangling or aliasing- which is important for generalization. Using context as the bonding element, we construct a pair of manifold untangling and tangling operators, known as tangling-untangling cycle (TUC). Untangling operator maps context-independent representations (CIR) in low-dimensional space to context-dependent representations (CDR) in high-dimensional space by inducing context as hidden variables. The tangling operator maps CDR back to CIR by a simple integral transformation for invariance and generalization. We also present the hierarchical extensions of TUC based on the Cartesian product and the fractal geometry. Despite the conceptual simplicity, TUC admits a biologically plausible and energy-efficient implementation based on the time-locking behavior of polychronization neural groups (PNG) and sleep-wake cycle (SWC). The TUC-based theory applies to the computational modeling of various cognitive functions by hippocampal-neocortical systems.
I Introduction
A key challenge in machine learning (ML) is balancing specification and generalization (a.k.a. over-fitting and under-fitting). We want a model that generalizes well (but not so simplistic that it underfits) and is well-specified to capture the essential patterns in the training data (but not so complex that it overfits). This subtle balance is often achieved through statistical techniques like cross-validation, regularization, and choosing the right model complexity. From a geometric perspective, the challenge is intrinsically related to the manifold structure of the data acquired from the physical world. To discover and exploit such manifold structure, nonlinearity is commonly believed to be essential, as witnessed by kernel trick in support vector machine [48] and nonlinear dimensionality reduction (e.g., IsoMAP [45] and locally linear embedding [40]). However, nonlinearity is often at odds with rigorous proof and the simplicity criterion in mathematical and scientific research.
In this paper, we draw inspiration from the discovery of place cells [34] and its related aliasing problem [29, 50]. Nature has discovered an elegant solution to efficient learning through the evolution of mammalian brains, but its secret has remained elusive. We shed some light on this greatest puzzle by revisiting the role of specification and generalization from a manifold untangling/tangling perspective [27]. Whitney’s embedding theorem in topology [31] has shown that it is easier to untangle a manifold in a high dimensional space. Such intuition inspires us to introduce context dependency as the clue for manifold untangling. Using class labels C as contextual variables, we can show an arbitrary manifold $P(\textbf{X})$ can be untangled (admitting linear dichotomy [6]) in the lifted space $P(\textbf{X},\textbf{C})$. The opposite direction (tangling operator) is a simple integral transform that collapses/aliases all variables sharing the same context. By associating tangling/untangling with context-independent representation (CIR) and context-dependent representation (CDR), we rigorously show how to strike an optimal tradeoff between specificity and generalization by a pair of tangling and untangling operators, which we call tangling-untangling cycle (TUC).
Using TUC as the building block, we can construct more sophisticated learning algorithms based on Cartesian products and fractal geometry [28]. It can be shown that the Cartesian product of TUC is still a TUC, which supports the multiscale extension from low-dimensional to high-dimensional space. To mirror the nested structure of data acquired from the physical world, fractal-inspired extension is constructed where the inter-scale context dependency is induced by an index variable (a.k.a. positioning trick [26]). The efficiency of the proposed learning algorithms is supported by their biological implementations based on polychronization neural groups (PNG) [24]. The implementation of TUC by sleep-wake cycle [13] finishes our journey at its starting point - the discovery of place cells in the hippocampus [34]. The presence and absence of external stimuli (i.e., default state vs. perturbation state) offer an energy-efficient and biologically plausible implementation of untangling and tangling operators in wake and sleep modes, respectively. Overall, the main contributions of this paper can be summarized as follows:
  • A geometric insight about the blessing and curse of dimensionality leads to a new understanding of specification and generalization in ML, which nicely fits the emerging direct-fit-to-nature paradigm.
  • A construction of manifold tangling-untangling cycle (TUC) for efficiency learning. Our construction represents a novel direct-fit-to-nature approach to unsupervised learning. Tangling and untangling operators are implemented by context dependency and integral transform, respectively.
  • A biological connection of TUC with sleep-wake cycle (SWC) and its related cognitive functions. Untangling transforms CIR to CDR for specification in the wake mode; and untangling maps CDR back to CIR for generalization in the sleep mode.
  • An application of TUC theory to model sensorimotor and social interactions. We use biological findings in various studies of cognitive functions as experimental support for the new theoretical framework.
II Geometric Intuition
Figure 1: Kernel trick achieves linear separability by mapping input data to a higher-dimensional feature space. $\phi:(x,y)\rightarrow(x,y,c)$, where $c=0,1$ denote red and blue classes, is a simple lifting idea that can be generalizable to a data manifold in arbitrary dimensions.
In graph theory, there is a well-known educational puzzle named “Three Utilities Problem”. This puzzle can be formalized as a problem in topological graph theory by asking whether the complete bipartite graph $K_{3,3}$, has a graph embedding in the plane. The impossibility of the puzzle corresponds to the fact that $K_{3,3}$ is not a planar graph. However, such a puzzle can be easily solved if the planarity constraint is removed. This is a simple example illustrating the blessing of dimensionality to the problem of manifold untangling. In topology, Whitney Embedding Theorem [31] states that any smooth manifold can be embedded into a Euclidean space of double its dimension. Based on this classical result, we can prove the following lemma:
Lemma (Linear separability in high-dimensional space)
Given two tangled manifolds $A,B\in R^{n}$, one can always linearly separate them in a high-dimensional space.
Proof. Using Whitney’s embedding theorem, we can also embed $A,B$ to a Euclidean space in $\tilde{A},\tilde{B}\in R^{2n}$. Then it is trivial to linearly separate $\tilde{A},\tilde{B}$ in $R^{4n}$.
In machine learning (ML), the kernel trick used by support vector machine [42] represents a good example for illustrating the blessing of dimensionality. As shown in Fig. 1, linearly non-separable data sets can become untangled (linearly separable) after being mapped to a high dimensional feature space. The construction or learning of features has been at the foundation of all existing models, including both biological and artificial neural networks. Recent direct-fit models suggest that we can solve the manifold untangling problem without feature engineering [20]. This view is also consistent with intelligence without representation [3] and supported by the discovery of concept cells [38, 37]. As summarized in [9], identity-related feature information of an object manifold can be simultaneously conveyed by the gnostic units in the IT cortex along with contextual information (e.g., position, pose, and scale), implying a high-dimensional latent space.
III Context Dependency and Integral Transform for Untangling and Tangling
Our intuition is based on observing the duality between specification (overfitting) and generalization (underfitting) in learning. An alternative approach to over/under-fitting is direct-fit-to-nature [20], which greatly inspired our approach in this work. Geometrically, we interpret overfitting/underfitting as manifold untangling and tangling, as shown in Fig. 2. Assuming sensory raw inputs are data belonging to a low-dimensional manifold embedded in a high-dimensional space, the objective of manifold untangling is to overfit the data for specification. This objective can be facilitated by embedding the manifold into a high-dimensional space, as Whitney’s embedding theorem suggests. Therefore, manifold untangling maps sensory raw input to a high-dimensional latent space where the fitting task becomes trivial (e.g., by linear dichotomy [6]). The dual process of manifold tangling is to underfit the data for generalization. A common example is aliasing [29], which reduces the dimensionality by collapsing the variables with a common state. We suggest an integral transform as a more general strategy than aliasing for generalization. In this section, we will study the pair of manifold untangling and tangling separately.
Figure 2: The tangling-untangling cycle (TUC) for solving overfitting/underfitting problems in efficient ML. Note its biological plausibility is consistent with sleep-wake cycle (SWC).
III-A Manifold Untangling: Dimensionality Expansion via Context Dependency
The intuition behind context dependency can be easily explained by the toy example of binary classification (e.g., cats vs. dogs). Suppose we collect sensory raw data associated with two classes $\textbf{X}_{0},\textbf{X}_{1}\in R^{n}$ corresponding to classes 0 and 1, respectively. The raw data in $R^{n}$ is tangled (e.g., blue vs. red dots as shown in Fig. 1) but they can be easily untangled by lifting the dimensionality by one - i.e., $(\textbf{X}_{c},c)\in R^{n+1}$ where $c=0,1$ (a linear classifier $c=1/2$ is sufficient). Such an idea of dimensionality expansion can be better articulated as a strategy of context dependency where the label is explicitly coupled with the training data as a context variable. Note that the introduction of label-based context dependency transforms supervised learning to unsupervised learning [2] by direct-fit-to-nature [20]. We can generalize the strategy of context dependency for high-dimensional labels/context as follows.
Proposition 1 (Specification). Context Dependency for Manifold Untangling.
Given sensory raw data $\textbf{X}\in R^{n}$ and its associated label $\textbf{C}\in R^{k}$, we can untangle the tangled data distribution $P(\textbf{X})$ by introducing context dependency as a joint PDF $P(\textbf{X},\textbf{C})\in R^{n+k}$.
The proof of proposition 1 is a straightforward extension of the strategy used in binary classification. By considering the discrete representation of label or context C, linear separability is guaranteed in the expanded higher-dimensional space.
[…]
III-B Manifold Tangling for Generalization: Dimensionality Reduction via Integral Transform
The dual operation of manifold tangling can be understood as the extension of aliasing [29]. The basic idea behind aliasing is to provide useful invariants by representing as equivalent states in which the same action/class is needed. Aliasing is good for generalization but has the risk of over-generalization. Such risk can be well managed by applying aliasing or its extended tangling operator to the untangled data in the high-dimensional space. The trick is to selectively remove or collapse hidden states where the context dependency dictates the criterion of selectivity. Using contextual information C as the guardrail, we are guaranteed a faithful generalization without the risk of over-generalization.
Proposition 2 (Generalization). Integral Transform for Manifold Tangling.
[…]
The pairing of manifold untangling and tangling operators forms a complete (TUC) cycle, as shown in Fig. 2. The connection of TUC with direct-fit models [20] is appealing - the objective of learning is not human-interpretable rules or world representations (i.e., “intelligence without representation” [3]). Instead, evolution has invented a localized computation for embodied cognition to interpolate over task-relevant manifolds (we will elaborate on this issue later when studying sensorimotor interaction). Manifold untangling offers a constructive solution to obtain over-parameterized models that are simple and parsimonious. More importantly, the connection of TUC with the evolutionary process is so powerful that we strongly believe it offers an important clue to unlock hidden secrets in developmental and ecological psychology. In summary, TUC forms the building block for our solution to efficient learning problems in general. Before we study the application of this new theory into neuroscience, we report several hierarchical extensions of TUC into chains and fundamental group of torus [21] to enrich our weaponry of constructing over-parameterized direct-fit models.
IV Hierarchical Extension and Connection with Fractal Geometry
First, we can spatially extend TUC with Cartesian products to form a fundamental group of the circles [21]. Such a dimensional increase is similar to the product PDF from the two marginals. Note that after manifold untangling, factorized linear dichotomies are conceptually similar to naive Bayes [39] (simple and parsimonious low-dimensional classifiers). Therefore, the Cartesian product preserves the independence of factorized distributions in the latent untangled space. In this manner, direct-fit models can be arbitrarily scaled to a high-dimensional space without suffering from the curse of dimensionality. Formally, we have the following result:
a)                                                      b)
Figure 3: Hierarchical extension of TUC: a) spatial extension by Cartesian product of TUC, which forms the fundamental group of the circles; b) temporal extension by alternating between untangling and tangling in time domain.
Proposition 3. (Cartesian Product of TUC forms the Fundamental Group of the Circle).
[…]
By recursively applying Cartesian products of TUC, we can construct a topological structure with increasing dimensionalities. Fig. 3a illustrates a toy example for such an extension of increasing dimensionalities (from a pair of circles to a torus). Analogous to the extension of circles to the fundamental group of circles, we can construct the fundamental group of TUC as $\textbf{S}^{1}\times...\times\textbf{S}^{1}$.
An alternative hierarchical extension of TUC is based on the temporal depth [35] where the process of untangling and tangling occurs in an alternating manner. Temporal depth or hierarchy is formed based on allostasis [35], as shown in Fig 3b. This construction will be useful for characterizing the dynamics of TUC, such as error feedback and temporal stability, especially in the presence of heterogeneous interactions (e.g., sensory interactions with motor). The alternating between tangling and untangling operators is also analogous to the alternating between sleep and wake modes in sleep-wake cycles, as we will elaborate in the next section.
Proposition 4. (Allostatic concatenation of TUCs forms temporal hierarchy.
Temporally hierarchical extension of TUCs can support continual learning by alternating between the processes of untangling and tangling in the time domain.
A more subtle construction of hierarchical TUC is based on fractal geometry [28]. We must extend the strategy context dependency across different scales to support fractal TUC. This can be implemented by the positioning trick [26], which was inspired by Hippocampal indexing theory [46]. The indexing variables $I$ induce inter-scale context dependency - e.g., in structure-mapping theory for analogy [16], higher-order relations are modeled based on the relation between objects instead of attributes of objects (so-called systematicity principle). The implementation of such principle requires the context to be the object X instead of class label C, which requires a fractal design such as the six-layered organization of cortical column. Such hierarchical organization of structures into TUCs might explain the magical number of six (e.g., six degrees of separation) or seven [30].
Figure 4: Illustration of sleep-wake cycle (SWC): untangling occurs during the wake time (perturbation state) and tangling occurs during the sleep time (default mode).
V Biological Connections with PNG and SWC
The biological implementation of the untangling operator is based on the connection of the polychronization neural groups (PNG) theory in [24] with the locally subspace untangling hypothesis in [9]. A direct consequence of direct-fit-to-nature is over-parameterization [20] to increase the predictive power without explicitly modeling the underlying generative structure of the world (i.e., intelligence without representation [3]). PNG offers a plausible computational model for supporting such mindless yet powerful generalizations. The number of PNG [24] can exponentially increase as more neurons are involved. Intuitively, one can argue that a high-level PNG with more neurons will have more representation power than a low-level PNG, which makes it a promising candidate for implementing the direct-fit-to-nature model [20].
We first extend Hebbian postulate [22] to make it work for PNG as follows - “A low-level PNG that $drives$ a high-level PNG to fire tends to wire together” [11]. Driving is hypothesized as a mechanism for generalizing cloning [50] to achieve specificity (e.g., unidirectional place cells [4] and context-dependent concept cells). The opposite direction from high-level to low-level generalizes aliasing [29] to achieve invariance or generalization (e.g., omnidirectional place cells and context-independent concept cells [38, 37]). Due to dimensionality expansion (the consequence of recursive untangling), the support of CDR would increase exponentially. PNG-based implementation can handle this combinatorial explosion because of its enormous capacity [24].
[…]
a)                                                      b)
Figure 5: Context dependency matters to both visual perception and motor control. a) PNG-based untangling for visual perception; b) PNG-based untangling for motor control.
The combination of PNG with SWC [13] allows us to study cognitive functions from a novel TUC perspective. Using TUC as the vocabulary, we attempt to construct a computational model for SWC as the first step toward a more comprehensive neural syntax [4]. As application examples, we consider three tasks here: spatial navigation, object recognition, and motor control.
Spatial Navigation. Navigation represents an easier learning problem that evolution solved long before the vision problem. Our understanding about this task has converged to a theory based on cognitive maps [47, 34, 50]. Cognitive maps organize knowledge to support flexible and adaptive behavior. How do mammalian brains encode 2D physical space information into 1D spike sequence? This question is at the core of understanding the relationship between space and time in its simplest form. The representation of spatial information (Euclidean or non-Euclidean) plays an essential role in the building of cognitive maps. A graph-based representation such as clone-structured causal graphs (CSCG) [17] has been widely adopted for encoding the physical space into action-related state space. During the construction of the CSCG model, the aliasing problem is solved by the cloning trick, which can be interpreted as a degenerated untangling operator.
Untangling during the wake mode (perturbation state) is demonstrated by place cells with unidirectional place fields. The 1D travel sequence dictates the spatio-temporal context, which helps resolve the ambiguity in space [4]. Tangling during the sleep mode (default state) is demonstrated by omnidirectional place cells [34] that do not require a temporal context. Memory consolidation [44, 49] implements the aliasing or tangling by binding all PNG which shared the same context C together. The collapse of cloned variables occurs at the hippocampus where the encoding of spatio-temporal context (index to the local map or position on the global map) is handled by CA3-CA1 collateral system (pp. 326 [4]). During the consolidation of semantic memory (collection of CIR) from episodic memory (collection of CDR), the PNG-based time-locking mechanism eliminates the spatio-temporal context to support context-independent inference for generalization.
[…]
Motor control. The duality between visual perception and motor control was observed in [36]. Based on this analogy, one can understand the untangling problem in the motor cortex similarly to that in the sensory cortex [51], as shown in Fig. 5b. The high-tangling state consists of options that are context-independent; while the low-tangling state is the collection of actions that are context-dependent. The importance of motor untangling to muscle-like actions/commands has been demonstrated in [41]. Sleep mode plays an important role in motor learning similarly to visual learning. During sleep, particularly during rapid eye movement (REM) and deep stages of non-REM sleep, the brain consolidates motor skills learned during the day. This process is essential for motor learning, helping to improve skills and refine motor tasks over time. Empirical evidence supporting this line of theoretical reasoning can be found in [7].
VI Computational Modeling of Sensorimotor and Social Interactions
TUC-based local computation for direct-fit inspired us to propose a new conceptual framework called “localized embodied cognition” with the locality principle in mind from problem formulation to solution algorithms. We still use the world “as its own model” [3], but the way local embodied cognition never attempts to solve a global problem locally. An organism needs to couple sensory with motor systems to achieve various goals but never stores knowledge about the global environment. Instead, the bodily interactions with the environment (situatedness) are persistently maintained by the neocortex (a single book)’s interaction with the hippocampus (a large library). The cognitive map in the hippocampus offers a situational context to interpret the sensorimotor interaction occurring in the neocortex. In this section, we study sensorimotor interaction and hippocampal-neocortical interaction using the newly-constructed neural syntax.
Figure 6: Sensorimotor cycle (SMC) consists of sensory untangling-tangling pair and motor untangling-tangling pair for predictive processing. Note that sensory and motor context-dependent representations (CDR) are task-relevant latent variables (the connection of chain); sensory and motor context-independent representation (CIR) are the interface with the external environment.
Neocortex as a Sensorimotor Machine.
One of the biggest mysteries in studying cortical columns is that they are structures without a function [23]. Instead, several canonical microcircuits such as Douglas-Martin model [10] and Haeusler-Maass model [18] have been studied for predictive coding in the literature. These models have revealed a remarkable correspondence between the microcircuitry of the cortical column and the connectivity required by predictive coding. Cortical columns do not have a function because they work like swiss-army knivies. By connecting with different sensory and motor areas, they offer a universal processing algorithm [32]. TUC-based modeling allows us to gain a deeper understanding of sensorimotor interaction at the system level without the distraction from lower-level implementations.
As shown in Fig. 6, the coupling of sensory and motor untangling-tangling pairs naturally forms a bigger cycle for the next level along the hierarchy of system neuroscience - modeling sensorimotor interaction by a pair of inverse and forward models [25]. Since more is different [1], the sensory motor cycle (SMC) is likely to be guided by a varying organizational principle, as initially demonstrated by [43]. Based on the duality between sensory and motor cortex [36], sensory CDR (after perceptual untangling) has to be linked with motor CDR (after motor untangling). They are task-relevant latent variables in the network dynamics of SMC [43]. We come up with the following conjecture but leave computational modeling for the mechanism of inverse and forward models in SMC and its relationship to SWC for a future study.
Conjecture (Context-dependent manifold untangling) The task of sensorimotor learning can be facilitated by simultaneously untangling the neural manifold of sensory and motor variables in the right positioning of latent space, where the position is specified by the index from the hippocampus as in Hippocampal indexing theory [46].
SMC-based Social Interaction
. Carrying the “more is different” principle one step further, we can envision a bigger cycle connecting a pair of SMCs. When one person becomes another’s environment, the communication between two SMCs forms another bigger cycle known as non-verbal communication. The communication bandwidth of such non-verbal communication is limited by the locality of our sensory organs. Even though within each neocortex, the hierarchical organization maximizes the information flow inward, social communication (flowing outward) has to wait for evolution to improve its efficiency.
Figure 7: The cycle connecting speech production by motor cortex and language comprehension by auditory cortex has a dual structure to SMC because social communication between two people is connected by CIR (tangled states).
The fundamental role played by natural language is to facilitate communication among people - from non-verbal to verbal. Fig. 7 shows a modification of SMC to abstract the verbal communication between two people A and B, where motor and auditory cortex are responsible for speech generation and language comprehension, respectively. Note that unlike the original SMC in Fig. 6, the interface between two TUCs is in the tangled state (CIR). This difference has a fundamental impact on our internal (self-conscious) and external (social) cognitive function. For internal communication between sensory and motor cortex, the interface (inverse and forward in Fig. 6) connects two untangled states and there is no bottleneck in terms of information flow. For external (social) communication between two different person, the interface (inverse and forward in Fig. 7) connects two tangled states and faces an information flow bottleneck.
According to the simulation theory of mind-reading [15], mirror neurons can help us understand and empathize with the emotions and intentions behind others’ non-verbal cues. Modeling the new cycle in Fig. 7for understanding the social nature of language might offer new insight to the neural exploitation hypothesis [14]. According to this hypothesis, social cognition and language are both connected with action through the support of mirror neurons. However, the difference between non-verbal and verbal communication seems to be determined by their organizational principles. Non-verbal communication such as sign language does not admit hierarchical organization; while verbal communication such as Chomsky’s hierarchy [5] in natural language offers an evolutionary solution to maximize the information flow within the new cycle.
VII Conclusions and Future Works
In this paper, we have constructed a new TUC-based theory to solve the long-standing learning problem. Exploiting the blessing of dimensionality by inducing context dependency allows us to untangle any data manifold so linear classifiers will be sufficient. Then the tangling operator based on integral transform collapses the induced context variables for generalization. Such mapping between tangled CIR and untangled CDR can be extended into hierarchical representations by Cartesian product and fractal geometry. The computational efficiency of TUC-based learning algorithms is supported by the biological implementation by SWC. Sleep and wake modes in SWC elegantly correspond to tangling and untangling operators due to the absence of sensory inputs during the sleep mode. SWC allows us to construct a neural syntax for an improved understanding of cognitive functions associated with sensorimotor interaction and social interaction.